# mystyle backgroundcolor=mygray, leftmargin=0.45em, rightmargin=0.45em, innertopmargin=10pt, innerbottommargin=10pt, skipabove=10pt, skipbelow=10pt

### v2

@@ Abstract
Software vulnerabilities pose significant risks to the security and integrity of software systems. Prior studies have proposed various approaches to vulnerability detection using deep learning or pre-trained models. However, there is still a lack of detailed explanations for understanding vulnerabilities beyond merely detecting their occurrence, which fails to truly help software developers understand and remediate the issues. Recently, large language models (LLMs) have demonstrated remarkable capabilities in comprehending complex contexts and generating content, presenting new opportunities for both detecting and explaining software vulnerabilities. In this paper, we conduct a comprehensive study to investigate the capabilities of LLMs in both detecting and explaining vulnerabilities, and we propose LLMVulExp, a framework that utilizes LLMs for these tasks. Under specialized fine-tuning for vulnerability explanation, our LLMVulExp not only detects the types of vulnerabilities in the code but also analyzes the code context to generate the cause, location, and repair suggestions for these vulnerabilities. These detailed explanations are crucial for helping developers quickly analyze and locate vulnerability issues, providing essential guidance and reference for effective remediation. We find that LLMVulExp can effectively enable the LLMs to perform vulnerability detection (e.g., achieving over a 90% F1 score on the SeVC dataset) and provide detailed explanations. We also explore the potential of using advanced strategies such as Chain-of-Thought (CoT) to guide the LLMs in concentrating on vulnerability-prone code, achieving promising results.
@@ I Introduction
A software vulnerability is a flaw or weakness in a system that can be exploited by an attacker to perform unauthorized actions [1, 2, 3, 4, 5]. These vulnerabilities can lead to severe consequences, including data breaches, financial losses, and damage to an organization’s reputation. The increasing complexity and interconnectedness of software systems introduce the great challenge of identifying and mitigating these vulnerabilities effectively.
@@
In the context of enhancing detection methods, the advent of Large Language Models (LLMs) offers a promising avenue. LLMs, with their advanced generative capabilities, have demonstrated potential in a wide range of applications, such as natural language processing and machine translation. These models can generate extensive textual content and provide contextually relevant information, making them natural candidates for tasks requiring detailed explanations. However, a substantial gap exists between the capabilities of LLMs and the specific requirements of vulnerability detection and explanation. They struggle to effectively detect vulnerabilities and provide accurate explanations without the domain knowledge of vulnerabilities, which requires a deep understanding of code structures, security contexts, and the intricate interplay between various software components. To bridge this gap, it is imperative to fine-tune LLMs specifically for vulnerability detection and explanation. This specialized training aims to endow LLMs with the necessary skills to accurately identify and explain software vulnerabilities.
Figure 1: Example of vulnerability detection result with and without explanation.
The challenges of this fine-tuning process are twofold: obtaining high-quality training data [1, 11, 12, 13, 14] and effectively adapting the models [15, 16, 17]. Curating relevant datasets that encompass diverse and representative software vulnerabilities is crucial. These datasets must capture various vulnerability types, from simple coding errors to complex logic flaws. Furthermore, fine-tuning requires sophisticated techniques to tailor LLMs to the nuanced requirements of vulnerability detection and explanation. Effective fine-tuning can enhance the model’s understanding of code structures and security contexts, thereby improving its detection and explanatory capabilities.
In this study, we explore the enhancement of LLMs for detecting and explaining software vulnerabilities by proposing an automated framework LLMVulExp, which involves leveraging prompt-based techniques to annotate explanatory information for open-source vulnerability data using advanced LLM models. The annotated explanatory information includes the location of the vulnerability and a detailed explanation. Following this, we employ Low-Rank Adaptation (LoRA) [16] fine-tuning methods to refine the LLMs, improving their detection capabilities and establishing a framework for evaluating their explanatory performance [18]. Additionally, we investigate the correlation between the capability of vulnerability detection and explanation, as well as utilize LLMs to follow a Chain-of-Thought (CoT) [19] strategy for concentrated vulnerability detection in key code snippets prone to software vulnerabilities. Furthermore, we address the evaluation of generated vulnerability explanations [20] by proposing new evaluation metrics and an automated review method using LLMs.
Through a comprehensive evaluation of two widely used vulnerability datasets (i.e., SeVC [21] and DiverseVul [11]), we find that LLMVulExp can enable the LLMs to detect vulnerabilities with a high accuracy (e.g., with an F1 score over 90% on SeVC) and provide an effective explanation. We also find that the key code extraction following a CoT strategy can improve the performance of vulnerability detection by a large margin.
@@
  • We present a comprehensive and effective workflow for training, inferring, and evaluating LLMs based on open-source vulnerability data.
  • We pioneer the exploration of the effectiveness of vulnerability explanation models and propose novel evaluation methods and dimensions for these explanations.
  • We analyze the interrelationship between explanatory and detection capabilities, offering valuable insights for the development of explainable vulnerability detection methodologies.
Paper Organization. Section II summarizes the related work. Section III presents the methodology of our study. Section IV discusses the results to our research questions. Section V discusses the implications of our study. Section VI discusses the threats to validity. Section VII concludes the paper.
@@ II Related Work
In this section, we summarize the related work on vulnerability detection and explanation, respectively.
@@ II-A Vulnerability Detection
Prior studies proposed a series of deep learning approaches [1, 2, 8, 9, 10] to detect vulnerabilities. These studies utilized labeled data with and without vulnerability to train neural networks and capture their semantic characteristics.
Apart from deep learning approaches, recent advancements in LLMs have significantly influenced the field of software vulnerability detection. Advanced methodologies such as zero-shot prompting [12, 22], in-context learning [23, 24], and fine-tuning [16, 15, 11] have been employed in vulnerability detection. Cheshkov et al. [20] evaluated the performance of the ChaptGPT and GPT-3 on vulnerability detection tasks and found they failed to classify vulnerable codes in binary and multi-label settings. Gao et al. [25] proposed a benchmark for using LLMs in vulnerability detection and validated that on simple datasets with few-shot prompting, LLMs can achieve performance that is comparable to or exceeds that of deep learning methods. Nong et al. [26] focused on evaluating the performance improvement of vulnerability detection tasks using chain-of-thought prompting. By introducing chain-of-thought forms based on the semantic structure of code, LLMs can achieve higher detection accuracy. Sun et al. [27] further assessed the true reasoning capabilities of LLMs by decoupling their vulnerability reasoning abilities. They found that supplementing LLMs with high-quality vulnerability-related knowledge and contextual information can enhance their performance. Yusuf et al. [28] discovered through experiments that natural language instructions enhance the performance of vulnerability detection tasks across multiple programming languages. Steenhoek et al. [29] surveyed and evaluated eleven LLMs for their capabilities of vulnerability detection, utilizing various types of prompts, including in-context learning and chain-of-thought, which indicates the accuracy limitation of directly applying LLMs to vulnerability detection without finetuning.
These studies underscore the challenges and opportunities of applying LLMs to vulnerability detection. Unlike prior methods, our research focuses on more practical and challenging explanation tasks, and we enhance the vulnerability understanding and analysis capabilities of LLMs through specialized fine-tuning.
@@ II-B Vulnerability Explanation
Although significant research progress has been made in applying deep learning techniques to vulnerability detection, effectively leveraging these techniques for vulnerability explanation remains a challenging issue. Currently, only a limited number of studies focus on the explanatory capabilities of deep learning based models regarding vulnerabilities.
The VulDeeLocator  [30] enhances a Bi-LSTM detector by incorporating an inner multiplication layer, which aids in forecasting vulnerable statements based on the outputs of this layer. In a similar vein, both IVDetect [2] and LineVul [31] develop vulnerability detection models that utilize subgraphs or attention weights derived from the trained detectors to identify vulnerable statements. VELVET [32] prioritizes vulnerable statements by integrating graph-based neural networks with sequence-based neural networks. LineVD [33] addresses statement-level vulnerability detection as a node classification challenge, employing graph neural networks combined with a transformer-based model on PDG for comprehensive learning. VulTeller [34] emphasizes control flows and taint flows to detect more accurate dependencies for localizing vulnerabilities. VulExplainer  [35] is proposed to locate the fine-grained information pertaining to the vulnerability based on graph neural networks. Coca [36] is a dual-perspective contrastive learning enhancement strategy to improve vulnerability explanation methods based on graph neural networks.
The semantic understanding of neural network offers the potential for vulnerability explanation models tailored to software development. However, accurate detection and effective explanation of vulnerabilities using LLMs remain substantial challenges for general-purpose code models. Our work explores the feasibility of fine-tuning specialized LLMs for vulnerability detection and explanation, addressing a crucial gap in this research domain.
@@ III Methodology
Figure 2: Overview of our study.
To address the current gap in generative vulnerability explanation models and enhance the ability of LLMs to detect and analyze software vulnerabilities, we propose a comprehensive framework for fine-tuning and evaluating specialized models for both vulnerability detection and explanation. Figure 2 presents an overview of our framework, namely, LLMVulExp. Specifically, our framework consists of four core stages: ❶ open-source vulnerability data collection, ❷ automated vulnerability explanation annotation based on prompt engineering, ❸ specialized fine-tuning of vulnerability detection and explanation through instruction-based fine-tuning, and ❹ evaluation of generative vulnerability explanation capabilities.
❶ Fine-tuning and Evaluation Vulnerability Dataset Collection: Enhancing the specialized capabilities of LLMs relies on large quantities of high-quality domain-specific data. In the context of vulnerability detection and explanation, the authenticity of the vulnerability code, the diversity of vulnerability types, and the sufficiency of examples of each type are particularly important. In this paper, we conduct the study on two datasets: (1) SeVC [21], which contains four core types and over 50,000 vulnerable code snippets, and (2) DiverseVul [11], which covers 295 real open-source projects and 150 CWE types.
@@
Semantics-based Vulnerability Candidate (SeVC) dataset includes 126 distinct Common Weakness Enumeration (CWE) types of vulnerabilities with 56,395 vulnerable samples and 364,232 non-vulnerable ones. The SeVC dataset is categorized into four primary groups based on the underlying causes of the vulnerabilities: Library/API Function Call, Array Usage, Pointer Usage and Arithmetic Expression. The SeVC dataset includes a significant number of vulnerabilities for each of its four types, making it suitable for fine-tuning and evaluating models designed to detect and explain specific types of vulnerabilities.
@@
DiverseVul is a C/C++ vulnerable source code dataset, which includes 18,945 vulnerable functions and 330,492 non-vulnerable functions derived from 7,514 commits, encompassing 150 CWEs. DiverseVul is currently the largest real-world C/C++ vulnerability dataset, characterized by longer code snippets, coverage of more diverse projects and CWE types, and lower label noise. Therefore, it is used to evaluate the capabilities of LLMs in handling more challenging real-world scenarios involving a broader range of vulnerability types. We select the top ten most frequent CWE types to construct a fine-tuning dataset, ensuring a sufficient number of samples and a well-defined number of classes for our multi-class detection task.
@@
We deduplicate the vulnerability samples using a hash method. Then, we downsample the non-vulnerability samples in a 1:1 ratio to the vulnerability samples to obtain a balanced dataset, aiming to reduce training overhead and avoid model bias. The dataset details are shown in Table 1. We split the processed dataset into training, validation, and test sets in an 80%: 10%: 10% ratio for both vulnerability and non-vulnerability samples to conduct our experiments.
@@
❷ Automated Vulnerability Explanation Annotation: Current open-source vulnerability datasets predominantly encompass information such as source code, vulnerability labels, CWE types, and commit messages. However, they lack detailed explanations of the vulnerability logic within the source code, presenting a significant challenge for vulnerability detection techniques to provide corresponding explanations for the detection results. Manually annotating real-world vulnerable code explanations requires extensive software development experience and in-depth knowledge of software vulnerabilities, which incurs high labor and time costs.
@@
To address this challenge, we propose an automated vulnerability explanation annotation method based on prompt engineering using LLMs. This method leverages the contextual learning and instruction-following capabilities of LLMs, utilizing prompt engineering to achieve large-scale, high-quality automated vulnerability explanation annotation. The prompts decompose the explanation goal into three sub-goals: vulnerability discrimination, code location, and specific explanation. By combining instruction-based prompt templates with well-annotated examples, we stimulate the model’s contextual learning capabilities, ensuring the effectiveness of vulnerability explanation annotation.
In this paper, we use GPT-3.5 [37], accessed via the API provided by OpenAI [38], to implement the annotation process. To address our research questions and experimental needs, we annotated 40,491 and 9,161 vulnerability explanation data points across two datasets, respectively. This effort fills the current gap in vulnerability explanation data.
@@
❸ Specialized Fine-Tuning of Vulnerability Detection and Explanation: The automated explanation annotation of open-source vulnerability data creates a large-scale dataset for vulnerability detection and explanation, addressing data bottleneck issues in the fine-tuning process. To enhance the vulnerability detection and explanation capabilities of LLMs (especially open-source models with lower computational overhead), we fine-tune the general LLMs to enable them to detect and explain specific types of vulnerabilities in real code. We use instruction-based prompts to guide tasks, helping LLMs correctly understand task goals and generate standardized outputs. To minimize the computational overhead of the fine-tuning process, we adopt the parameter-efficient fine-tuning technique LoRA [16], significantly reducing time and space costs.
@@
❹ Evaluation of Generative Vulnerability Explanation: The current lack of research on model-generated vulnerability explanations highlights the need for effective evaluation methods for LLM-generated explanations. Similar to the challenges in annotation, manual evaluation demands significant human and time resources. Additionally, the effectiveness of vulnerability explanations must be assessed across multiple dimensions to ensure they provide valuable assistance to developers in real-world scenarios. To efficiently evaluate vulnerability explanations, we propose an evaluation method based on three dimensions: accuracy, clarity, and actionability. By leveraging prompt engineering, we develop an automated LLM evaluation method. Expert manual verification is employed to validate both the quality of the specialized vulnerability explanation model’s outputs and the feasibility of the LLM-based automated evaluation scheme.

### v3

@@ Abstract
Software vulnerabilities pose significant risks to the security and integrity of software systems. Although prior studies have explored vulnerability detection using deep learning and pre-trained models, these approaches often fail to provide the detailed explanations necessary for developers to understand and remediate vulnerabilities effectively. The advent of large language models (LLMs) has introduced transformative potential due to their advanced generative capabilities and ability to comprehend complex contexts, offering new possibilities for addressing these challenges. In this paper, we propose LLMVulExp, an automated framework designed to specialize LLMs for the dual tasks of vulnerability detection and explanation. To address the challenges of acquiring high-quality annotated data and injecting domain-specific knowledge, LLMVulExp leverages prompt-based techniques for annotating vulnerability explanations and fine-tunes LLMs using instruction tuning with Low-Rank Adaptation (LoRA), enabling LLMVulExp to detect vulnerability types in code while generating detailed explanations, including the cause, location, and repair suggestions. Additionally, we employ a Chain-of-Thought (CoT) based key code extraction strategy to focus LLMs on analyzing vulnerability-prone code, further enhancing detection accuracy and explanatory depth. Our experimental results demonstrate that LLMVulExp achieves over a 90% F1 score on the SeVC dataset, effectively combining high detection accuracy with actionable and coherent explanations. This study highlights the feasibility of utilizing LLMs for real-world vulnerability detection and explanation tasks, providing critical insights into their adaptation and application in software security.
@@ I Introduction
A software vulnerability is a flaw or weakness in a system that can be exploited by an attacker to perform unauthorized actions [1, 2]. These vulnerabilities can lead to severe consequences, including data breaches, financial losses, and damage to an organization’s reputation. The increasing complexity and interconnectedness of software systems introduce the great challenge of identifying and mitigating these vulnerabilities effectively.
@@
In the context of enhancing detection methods, the advent of Large Language Models (LLMs) [8, 9, 10] offers a promising avenue. With their advanced generative capabilities, LLMs have demonstrated significant potential across a wide range of applications, such as natural language processing and machine translation. These models can generate extensive textual content and provide contextually relevant information, making them well-suited for tasks that require detailed explanations. However, a considerable gap exists between the current capabilities of LLMs, particularly open-source models, and the specific requirements of vulnerability detection and explanation. Without domain-specific knowledge of vulnerabilities, LLMs struggle to effectively detect vulnerabilities and offer accurate explanations, which demand a deep understanding of code structures, security contexts, and the intricate interplay between various software components. To bridge this gap, it is essential to enhance LLMs with specialized capabilities for vulnerability detection and explanation. Fine-tuning open-source LLMs [11, 12, 13] on task-specific data presents an effective path toward achieving model specialization.
Figure 1: Illustration of Vulnerability Detection Explanation.
The main challenges in this specialization fine-tuning lie in two areas: (1) acquiring high-quality vulnerability explanation data [1, 14, 10, 15, 16], and (2) injecting vulnerability-related domain knowledge into LLMs [17]. The improvement of model specialization depends heavily on high-quality, domain-specific data. However, in the field of vulnerability detection, there is a significant lack of annotated explanation data. Additionally, annotating such data typically requires costly manual labeling by domain experts, making the collection of large-scale annotated datasets a major challenge for model fine-tuning. Additionally, due to the inherent complexity of vulnerability detection and explanation tasks, merely providing task descriptions in prompts is insufficient for generating accurate answers. The model requires additional, critical vulnerability domain knowledge to identify and explain vulnerabilities accurately.
To tackle these challenges, this study introduces LLMVulExp, an innovative automated framework designed to specialize open-source LLMs for software vulnerability tasks. Our framework directly addresses the two key challenges identified: the acquisition of high-quality annotated explanatory data and the effective injection of domain-specific vulnerability knowledge into LLMs. To address the first challenge, we leverage prompt-based techniques to automatically annotate explanatory information for vulnerability data. This annotated data includes both the vulnerability location and a detailed explanation, providing a rich dataset for fine-tuning. This approach reduces reliance on costly manual labeling by domain experts, enabling the collection of large-scale, annotated datasets necessary for model specialization. In addition, we propose new evaluation metrics for vulnerability explanations and an automated ’LLM-as-a-judge’ evaluation method [18, 19], ensuring a robust and scalable approach to verify the correctness and coherence of the generated explanations. For the second challenge, we apply instruction tuning [20] with Low-Rank Adaptation (LoRA) fine-tuning methods [12] to refine the open-source LLMs for injecting sufficient vulnerability patterns and knowledge. By focusing on task-specific adaptation, we guide the model to develop a deeper understanding of vulnerability semantics. To further enhance the model’s structural understanding of vulnerabilities, we adopt a Chain-of-Thought (CoT) strategy [21], guiding the model to concentrate on key code snippets most prone to vulnerabilities.
@@
  • We introduce a comprehensive and effective workflow for training, inferring, and evaluating LLMs specifically on vulnerability detection and explanation tasks.
  • We pioneer the exploration of the effectiveness of LLM-based vulnerability explanation generation and propose a tailored evaluation method and metrics for these explanations.
  • We conduct extensive experiments to evaluate and analyze the feasibility of LLM-based vulnerability explanations, providing valuable insights for the practical application of specialized models in vulnerability detection.
Paper Organization. Section II summarizes the related work. Section III presents the methodology of our study. Section IV discusses the results of our research questions. Section V discusses the implications of our study. Section VI discusses the threats to validity. Section VII concludes the paper.
@@ II Related Work
In this section, we review the related work in two key areas: vulnerability detection and explanation.
@@ II-A Vulnerability Detection
Prior studies proposed a series of deep learning approaches [1, 2, 5, 6, 7] to detect vulnerabilities. These methods have employed labeled vulnerability data to train neural networks, enabling the models to capture semantic features associated with vulnerabilities.
The advent of LLMs has significantly influenced the field of vulnerability detection. Techniques such as zero-shot prompting [10, 9], in-context learning [22, 23], and fine-tuning [12, 11, 14] have been explored to enhance LLM-based vulnerability detection. Cheshkov et al.[18] evaluated ChatGPT and GPT-3, highlighting their inability to classify vulnerable code accurately in binary and multi-label settings. Gao et al.[24] proposed a benchmark for LLMs in vulnerability detection, demonstrating that with few-shot prompting on simpler datasets, LLMs can perform comparably to deep learning-based methods. Nong et al.[25] showed that chain-of-thought prompting, based on the semantic structure of code, improves detection accuracy. Sun et al.[26] found that supplementing LLMs with high-quality vulnerability-related knowledge enhances their performance. Yusuf et al.[27] observed that natural language instructions boost vulnerability detection across multiple programming languages. Steenhoek et al.[28] surveyed eleven LLMs, indicating the limitations of applying LLMs directly to vulnerability detection without fine-tuning.
These studies underscore the challenges and potential of using LLMs for vulnerability detection. Unlike these methods, our research focuses on more practical and challenging explanation tasks, and we enhance the vulnerability understanding and analysis capabilities of LLMs through specialized fine-tuning.
@@ II-B Vulnerability Explanation
Although substantial progress has been made in applying deep learning techniques to vulnerability detection, effectively leveraging these methods for vulnerability explanation remains a significant challenge. Despite the critical importance of reducing reliance on security experts and aiding developers in timely and thorough vulnerability mitigation, this area remains underexplored. Only a few studies focus on the explanatory capabilities of deep learning based models for vulnerabilities.
VulDeeLocator [29] enhances a Bi-LSTM detector with an inner multiplication layer to forecast vulnerable statements. IVDetect [2] and LineVul [30] utilize subgraphs or attention weights from trained detectors to identify vulnerabilities. VELVET [31] integrates graph-based and sequence-based neural networks to prioritize vulnerable statements. LineVD [32] treats statement-level vulnerability detection as a node classification task, combining graph neural networks with a transformer-based model on the Program Dependency Graph(PDG). VulTeller [33] focuses on control and taint flows to detect dependencies for localizing vulnerabilities. VulExplainer [17] uses graph neural networks to locate fine-grained vulnerability information.
While neural network-based models offer promising semantic understanding for vulnerability explanation, accurately detecting and explaining vulnerabilities using LLMs remains a substantial challenge for general-purpose code models. Our work explores the feasibility of fine-tuning specialized LLMs for both vulnerability detection and explanation, addressing a critical gap in the research literature.
@@ III Methodology
Figure 2: Overview of the LLMVulExp Framework for Explainable Vulnerability Detection via LLM Fine-Tuning.
To address the current gap in generative vulnerability explanation models and enhance the ability of LLMs to detect and explain software vulnerabilities, we propose a comprehensive framework for fine-tuning and evaluating specialized LLMs for both vulnerability detection and explanation tasks. Figure 2 presents an overview of our framework, namely, LLMVulExp. Specifically, our framework consists of four core stages: ❶ fine-tuning and evaluation vulnerability dataset collection, ❷ automated vulnerability explanation annotation based on prompt engineering, ❸ specialized fine-tuning of vulnerability detection and explanation through instruction-based fine-tuning, and ❹ evaluation of generative vulnerability explanation capabilities.
❶ Fine-tuning and Evaluation Vulnerability Dataset Collection: Enhancing the specialized capabilities of LLMs requires large quantities of high-quality domain-specific data. In the context of vulnerability detection and explanation, it is crucial to ensure the authenticity of the vulnerability code, the diversity of vulnerability types, and the sufficiency of examples for each type. In this paper, we conduct the study on two C/C++ function-level vulnerability datasets: (1) SeVC [34], which contains four core vulnerability types and over 50,000 vulnerable code snippets, and (2) DiverseVul [14], which covers 295 real open-source projects and 150 CWE types. We select C/C++ as the programming language due to its diverse range of vulnerability types and examples, as well as the substantial dependence of fine-tuning on the availability of large-scale data.
@@
Semantics-based Vulnerability Candidate (SeVC) dataset includes 126 distinct Common Weakness Enumeration (CWE) types, comprising 56,395 vulnerable samples and 364,232 non-vulnerable ones. The SeVC dataset is categorized into four primary vulnerability classes based on the underlying causes: Library/API Function Call, Array Usage, Pointer Usage and Arithmetic Expression. Each category contains a substantial number of samples, making SeVC ideal for fine-tuning and evaluating models designed to detect and explain specific vulnerabilities. Therefore, we conduct both binary and multi-class detection tasks on the SeVC dataset.
@@
DiverseVul is a comprehensive C/C++ vulnerability dataset that includes 18,945 vulnerable functions and 330,492 non-vulnerable functions, extracted from 7,514 commits, covering 150 CWEs. As the largest real-world C/C++ vulnerability dataset, DiverseVul is characterized by longer code snippets, a broader range of projects, more diverse vulnerability types, and lower label noise. It serves as a benchmark for evaluating the capability of LLMs in handling real-world scenarios with a wider variety of vulnerabilities. For our fine-tuning and evaluation target, we selected the top ten most frequent CWE types, ensuring a sufficient number of samples and a well-defined number of classes for our multi-class detection task.
@@
To ensure data quality, we deduplicate the vulnerability samples using the SHA-256 [35] hash method and downsample the non-vulnerable samples at a 1:1 ratio to match the vulnerable samples, achieving a balanced dataset, which aims to reduce training overhead and mitigates potential model bias. The details of the datasets are shown in Table I. We split the processed dataset into training, validation, and test sets in an 80%: 10%: 10% ratio for both vulnerable and non-vulnerable samples to conduct our experiments.
@@
❷ Automated Vulnerability Explanation Annotation: Traditional open-source vulnerability datasets mainly contain source code, vulnerability labels, CWE types, and commit messages. However, they often lack detailed explanations of the vulnerability logic within the source code, posing a significant challenge for vulnerability detection techniques that aim to provide meaningful explanations for detected issues. Manually annotating real-world vulnerable code explanations requires extensive software development experience and a deep understanding of software vulnerabilities, which incurs high labor and time costs [36]. As the scale of manually annotated data remains a significant challenge, an increasing number of researchers are leveraging the powerful generative capabilities of LLMs to synthesize data. This approach has been validated in various domains through improvements in downstream task performance metrics. However, in the specialized field of vulnerability detection, where domain expertise is critical, the feasibility of synthesizing data has yet to be fully validated.
@@
To address this challenge, we introduce an innovative automated vulnerability explanation annotation method based on prompt engineering with LLMs. This method capitalizes on the contextual learning and instruction-following capabilities of LLMs, enabling large-scale, high-quality automated synthesis of vulnerability explanations. Our approach decomposes the explanation task into three sub-goals: (1) vulnerability discrimination, (2) identifying the location of the vulnerability in the code, and (3) providing a specific explanation. The model is guided by instruction-based prompt templates, paired with well-annotated examples, which stimulate its contextual learning capabilities, ensuring that the generated explanations are both accurate and informative.
We implement this annotation process using GPT-3.5 [37], accessed via the API provided by OpenAI [38] due to its balance of annotation cost and efficiency. As part of our experimental study, we annotated 40,491 and 9,161 vulnerability explanation data points across two datasets, respectively. This effort fills a significant gap in the availability of vulnerability explanation annotations, enabling more effective training and evaluation of vulnerability detection and explanation models. To validate the effectiveness of the synthesized data, we indirectly evaluated the performance of the fine-tuned model in vulnerability detection tasks across multiple scenarios. Additionally, we propose evaluation metrics for vulnerability explanation quality, along with both manual and automated evaluation methods, to directly assess the quality of the generated explanations.
@@
❸ Specialized Fine-Tuning of Vulnerability Detection and Explanation: The automated annotation of vulnerability explanations for open-source data creates a large-scale dataset, effectively addressing data bottlenecks in the fine-tuning process for vulnerability detection and explanation. To enhance the detection and explanation capabilities of LLMs (especially open-source models with lower computational overhead), we fine-tune general LLMs to specialize in detecting and explaining specific types of vulnerabilities in real-world code. Instruction-based prompts are used to guide these tasks, helping LLMs to accurately understand the objectives and generate standardized outputs. To minimize the computational cost of the fine-tuning process, we employ the parameter-efficient fine-tuning method LoRA [12], which significantly reduces both time and space requirements.
@@
❹ Evaluation of Generative Vulnerability Explanation: The limited research on model-generated vulnerability explanations underscores the need for robust evaluation methods tailored to LLM-generated outputs. Similar to annotation challenges, manual evaluation requires considerable human and time resources. Moreover, effective vulnerability explanations must be assessed across multiple dimensions to ensure they offer practical value to developers in real-world scenarios. To address this, we propose an evaluation framework based on three critical dimensions: accuracy, clarity, and actionability. These dimensions collectively measure the correctness, comprehensibility, and practical applicability of the generated explanations. To enhance evaluation efficiency, we introduce an automated evaluation method based on LLMs. This method leverages prompt engineering to enhance evaluation efficiency while ensuring reliable assessments. Additionally, expert manual verification is employed to validate both the quality of the outputs from the specialized vulnerability explanation model and the feasibility of the LLM-based automated evaluation scheme.